# Power allocation for cell-free MIMO integrated sensing and communication

### v1

@@ VII-A Convergence analysis
Proof. The proof is given in Appendix A. $\blacksquare$
Definition. A function $L({\bm{\rho}})$ is Lipschitz gradient continuous over a set $\Omega$, if there exists a constant $C>0$, $\|\nabla L({\bm{\rho}}_{i})-\nabla L({\bm{\rho}}_{j})\|\leq C\|{\bm{\rho}}_{i}-{\bm{\rho}}_{j}\|$ for any ${\bm{\rho}}_{i},{\bm{\rho}}_{j}\in\Omega$ [58, 59].
Theorem 1. Given a nonempty convex set $\Omega$ where the objective function $L({\bm{\rho}}_{i})$ is bounded over its interior points, if the objective function is Lipschitz gradient continuous over $\Omega$, the ILS-based descent direction iterative algorithm will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$. In contrast, if it is not Lipschitz gradient continuous, the sequence $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ may converge to a Lipschitz gradient noncontinuous point where the norm of the Hessian matrix may approach infinity.
@@
Theorem 1 provides the convergence analysis for the proposed PP-MCG-ILS and PP-MSD-ILS algorithms.
@@
The feasible domain $\Omega\subset\mathbb{R}^{N}$ is closed, bounded, and convex. All CRLB-involved matrices, e.g., ${\bm{P}}(\rho)$, ${\bm{Y}}(\rho)$, and their inverses are uniformly bounded and positive definite over the set $\Omega$.
Firstly, each term $f_{\min,n}({\bm{\rho}})$ and $f_{\max,n}({\bm{\rho}})$ is piecewise quadratic, and its gradient is Lipschitz continuous due to continuity and boundedness. So does ${f}_{\rm s}({\bm{\rho}})$. Then, since ${\bm{C}}_{L}({\bm{\rho}})$ is uniformly positive definite and smooth in ${\bm{\rho}}$, it follows that: $f_{l}({\bm{\rho}})$ is continuously differentiable, and so is
@@
Given that $\nabla f_{l}({\bm{\rho}})$ is smooth and bounded (due to the matrix inverse being smooth on positive definite matrices), $\nabla\tilde{f}_{l}({\bm{\rho}})$ is Lipschitz continuous. A similar argument applies to $\tilde{f}_{v}({\bm{\rho}})$. Hence, $\nabla_{\bm{\rho}}\alpha({\bm{\rho}})$ and $\nabla_{\bm{\rho}}\alpha_{\rm s}({\bm{\rho}})$ are Lipschitz continuous over $\Omega$, and so are $\nabla_{\bm{\rho}}L({\bm{\rho}})$ and $\nabla_{\bm{\rho}}L_{\rm s}({\bm{\rho}})$.
Additionally, since ${\bm{\Theta}}$ is a constant matrix with $\|{\bm{\Theta}}\|_{2}=1$, and $\nabla_{\bm{\rho}}f({\bm{\rho}})$ is Lipschitz continuous, the composition $\nabla\mathcal{L}({\bm{\varpi}})={\bm{\Theta}}\nabla L({\bm{\rho}})$ is also Lipschitz continuous. There exists a constant $C>0$ such that $\forall{\bm{\varpi}}_{1},{\bm{\varpi}}_{2}\in\Omega_{\bm{\varpi}}$
$\|\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{1})-\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{2})\|_{2}\leq C\|{\bm{\varpi}}_{1}-{\bm{\varpi}}_{2}\|_{2}.$
@@
VII-B Complexity analysis
We observe that the additional computational complexity of the PP-MCG-ILS algorithm, compared to the PP-MSD-ILS algorithm, primarily stems from the computation of $\varsigma_{i}^{\rm b}$ and $\varsigma_{i}^{\rm HS}$ at each iteration, which requires approximately $4N$ multiplications and 2 divisions. In both algorithms, the dominant computational burden lies in evaluating the objective function during the ILS and computing the gradients for direction updates, which are inherently tied to the CRLB expressions. Therefore, a more meaningful comparison of total computational cost focuses on the number of ILS steps per iteration. This number is influenced by the step size threshold, given by the right-hand side of (67), the initial step size $\upsilon_{0,i}$ in (52), and the scaling factor. At a given iterate ${\bm{\rho}}_{i}$, a larger descent direction norm $\|{\bm{d}}_{i}\|$ generally leads to a smaller $\upsilon_{0,i}$ and a reduced step size threshold. As a result, the number of ILS steps is expected to be comparable between the PP-MCG-ILS and PP-MSD-ILS algorithms. Given this, the primary metrics for evaluating the relative performance of these two algorithms become their convergence rate and the quality of their steady-state solutions.
We now qualitatively compare the computational complexity between algorithms employing ILS and those without ILS. As previously discussed, the number of ILS steps significantly impacts the number of evaluations of the objective function. Thus, the computational complexity of the benchmark algorithms (without ILS) can be considerably lower at each iteration during step 1 compared to the PP-MCG-ILS and PP-MSD-ILS algorithms. However, all four algorithms have similar complexity in step 4 due to the inherently intensive gradient calculations involved. Moreover, the overall computational complexity heavily depends on the convergence rates, or equivalently, the total number of iterations required for convergence. As will be demonstrated in the subsequent simulation results, the proposed PP-MCG-ILS algorithm exhibits a notably higher convergence rate compared to its counterparts, with or without ILS. This superior convergence performance significantly reduces the overall computational effort, particularly benefiting the demanding computations of gradients.
VIII Simulation Results
In this section, we conduct some simulations to evaluate the proposed PP-MCG-ILS and PP-MSD-ILS algorithms for PA optimization. All APs, including $N$ single antenna transmitters and $K$ single antenna receivers, are assumed to be stationary. We employ the orthogonal chirp division multiplexing (OCDM) [60] waveform with a single Gaussian pulse shaping. The carrier frequency is $3$GHz. The lowpass equivalent is formulated by,
@@ VIII Simulation Results
VIII-A CRLBs
We first verify the CRLB in comparison with the MLE results. To maintain consistency with existing works [38, 40, 41, 39, 36, 34, 35], we use the signal energy-to-noise ratio (SENR), defined as $\tilde{\delta}_{\rm cn}\triangleq P/\sigma_{\rm cn}^{2}$, instead of the SCNR $\delta_{\rm cn}$ in evaluating the CRLBs. Due to the high complexity of four-dimensional MLE, we consider a simple $4\times 3$ nonsymmetric multi-static radar configuration given in Fig. 2. Consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=1$ kHz. The velocity of the target is $(20,30)$ m/s.
@@ VIII-A CRLBs
VIII-B Power Allocation for ISAC
The performance of the proposed PA algorithms is evaluated by comparison with the PP-NCG and PP-NSD benchmark algorithms. To ensure communication quality, a larger number of APs are designated for transmitting ISAC signals, while a smaller subset is used for receiving target-reflected signals. Without loss of generality, we consider a CF MIMO-ISAC system comprising 10 transmitters and 2 receivers, as illustrated in Fig. 2. To guarantee the CRLB effectiveness, we assume the minimum PA coefficient vector is ${\bm{\rho}}_{\min}=0.01\times{\bm{1}}_{N}$. The PA budget is ${\bm{\rho}}_{\max}=0.3\times{\bm{1}}_{N}$. Consider a target moving slowly with velocity $l_{v}=[4,5]m/s$. A typical modulus squared RCS matrix is given by $\begin{bmatrix}0.37&0.70&1.38&0.65&2.40&0.18&0.25&0.82&0.14&0.35\\ 0.40&2.05&1.14&0.42&0.03&0.11&2.02&1.65&1.24&3.65\end{bmatrix}^{\rm T}$. The channel gains between the user and the transmitters are assumed to be $[2.11,12.57,5.63,0.75,0.61,1.75,0.20,2.34,14.79,9.68]^{\rm T}$. In this ISAC setup, the CRLB thresholds for the position and velocity are given by $[250,0.13]$.
@@ VIII-B Power Allocation for ISAC
The PA vector is initialized as the normalized PA vector ${\bm{\rho}_{1}}={\bm{1}}_{N}/N$, unless noted otherwise. The initial penalty factor is set to be $\mu_{1}=10^{4}$. We hereafter consider the ILS controlling factor $\epsilon_{L}=1e-3$, the initial step size controlling factor $\epsilon_{\upsilon}=0.9$, the penalty factor scaling coefficient $\phi=10$, and the penalty threshold $\epsilon_{\mu}=10^{-3}$, unless noted otherwise. The other simulation conditions are given in Table I. Note that the presented SINRs, PA vectors, and CRLBs, are obtained by averaging 100 points after convergence.
Firstly, we evaluate the convergence and steady-state performance of the proposed algorithms. The SINR learning curves of four algorithms are illustrated in Fig. 4. As apparently seen in Fig. 4, the PP-MCG-ILS has the highest convergence rate, followed by the PP-MSD-ILS, while the counterparts without ILS presents comparatively slow convergence. Additionally, the benchmark algorithms without ILS exhibit numerical fluctuations, particularly under large fixed step sizes. This behavior suggests potential instability and divergence risks. Consequently, we refrain from increasing the step sizes further during simulations to avoid compromising convergence reliability.
The evolution of the PA coefficients and corresponding CRLB values with respect to the number of iterations is illustrated in Fig. 5 and Fig. 6, respectively. The results show that the PP-MCG-ILS algorithm achieves significantly faster convergence compared to the PP-NCG algorithm. Ultimately, all four algorithms converge to a common PA vector, given by ${\bm{\rho}}=[0.01,0.30,0.0638,0.01,\dots,0.01,0.30,0.2762]^{\rm T}$, which satisfies the sensing QoS constraints. Notably, more power is allocated to transmitters 2, 9, and 11 due to their higher communication channel gains. Furthermore, the steady-state CRLB values remain within the predefined threshold limits, confirming the validity and effectiveness of the proposed optimization framework.
@@
Figure 4: SINR with sequential penalty factor
@@
The proposed sequential penalty-based algorithms leverage the optimal solution from the previous iteration loop as the initial point for the current loop under an updated penalty factor. To further explore the impact of initialization, we compare the performance of different initial PA strategies under a fixed penalty factor. In addition to the commonly used uniform PA initialization, we introduce a heuristic PA scheme, where the initial allocation is proportional to the communication channel gain, given by ${\bm{\rho}}_{1}={\bm{g}}/\sum_{n}|g_{n}|^{2}$. The results are illustrated in Figs. 7 and 8. When comparing to sequential penalty-based counterparts in Fig. 4, the PP-MCG-ILS and PP-NCG algorithms initialized with uniform PA exhibit similar convergence behavior. However, the PP-MSD-ILS and PP-NSD algorithms show degraded convergence, highlighting the superior convergence properties of CG-based algorithms over SD-based methods under identical conditions. When employing the heuristic PA initialization, all algorithms demonstrate significantly improved convergence rates compared to the uniform PA case. Moreover, both PP-NCG and PP-NSD achieve noticeable performance gains relative to their sequential penalty-based counterparts, particularly PP-NCG. In contrast, the PP-MCG-ILS algorithm shows robust and superior convergence behavior, and PP-MSD-ILS exhibits a slight degradation.
@@
We now investigate the impact of the CRLB thresholds on the performance of the proposed PP-MCG-ILS algorithm. As shown in Fig. 9, relaxing the location CRLB thresholds leads to a gradual increase in the steady-state SINR, which eventually stabilizes. This trend aligns with the behavior observed in Fig. 9, where the achieved location CRLB value closely approaches its threshold and then converge. Furthermore, it is noteworthy that the velocity CRLB remains largely unaffected by variations in the location CRLB thresholds. In contrast, Fig. 10 shows that the performance of the PP-MCG-ILS algorithm exhibits negligible variation with the increased velocity CRLB threshold. The observed CRLB saturation with respect to the corresponding CRLB threshold may be an inherent outcome of the introduced power budget constraints.
@@
VIII-C Power Allocation for pure sensing
@@
We now simply verify the effectiveness of the proposed P-MCG-ILS algorithm for total power minimization in joint location and velocity estimation. The initial penalty factor is set to be $\mu_{1}=1$, $\sigma_{\upsilon}=0.2$ and the penalty threshold $\epsilon_{\mu}=10^{-6}$. The other simulation conditions are the same as the PP-MCG-ILS algorithm in Fig. 4.
@@
The learning results for total power optimization and the corresponding CRLB performance are illustrated in Fig. 11. As seen in Fig. 11, the algorithm demonstrates rapid convergence. Under the same CRLB threshold settings, the total power consumption in the pure sensing scenario is significantly lower than that in the ISAC case. This is expected, as ISAC systems must allocate sufficient power to meet both sensing and communication requirements, particularly for maximizing communication performance. Moreover, both the location and velocity CRLBs in the pure sensing case closely align with their respective thresholds. In contrast, in the ISAC scenario (refer to Fig. 6), the velocity CRLB matches its threshold loosely. This difference arises because, in pure sensing, any enhancement in CRLB performance directly benefits from increased sensing power. However, in ISAC, power allocation is optimized primarily for communication performance, with CRLB constraints being satisfied, potentially loosely, within acceptable margins.
IX Conclusions

### v2

@@ VI-A Convergence Analysis
Definition. A function $L({\bm{\rho}})$ is Lipschitz gradient continuous over a set $\Omega$, if there exists a constant $C>0$, $\|\nabla L({\bm{\rho}}_{i})-\nabla L({\bm{\rho}}_{j})\|\leq C\|{\bm{\rho}}_{i}-{\bm{\rho}}_{j}\|$ for any ${\bm{\rho}}_{i},{\bm{\rho}}_{j}\in\Omega$ [43].
Theorem 1. Given a nonempty convex set $\Omega$ with the objective function $L({\bm{\rho}}_{i})$ bounded over its interior points, if the objective function is Lipschitz gradient continuous over $\Omega$, the ILS-based iterative algorithm will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$. In contrast, if it is Lipschitz gradient non-continuous, sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ may converge to a Lipschitz gradient noncontinuous point with the norm of the Hessian matrix approaching infinity.
Proof. The proof is given in Appendix C. $\blacksquare$
Theorem 1 provides the convergence analysis for the proposed iterative algorithms.
@@
Firstly, all CRLB-involved matrices, e.g., ${\bm{P}}(\rho)$, ${\bm{Y}}(\rho)$, and their inverses are smooth on ${\bm{\rho}}$ over the feasible domain $\Omega\subset\mathbb{R}^{N}$. Moreover, ${\bm{C}}_{L}({\bm{\rho}})$ is uniformly positive definite. It follows that ${\bm{C}}_{L}({\bm{\rho}})$ is continuously differentiable, and so is ${\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\}$. By the chain rule, we have
@@
where $\nabla f_{l}({\bm{\rho}})=-{\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\nabla({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T}){\bm{C}}_{L}({\bm{\rho}})\}$. Since the matrix inverse and trace are smooth on the set of positive definite matrices, $\nabla f_{l}({\bm{\rho}})$ is continuously differentiable. Moreover, all matrices involved (including ${\bm{C}}_{L}({\bm{\rho}})$) have bounded spectral norm, so the operator norm of the Hessian of $f_{l}({\bm{\rho}})$ is also bounded. Consequently, $\nabla\tilde{f}_{l}({\bm{\rho}})$ is Lipschitz continuous. A similar argument applies to $\nabla\tilde{f}_{v}({\bm{\rho}})$. On the other hand, $f_{\min,n}({\bm{\rho}})$, $f_{\max,n}({\bm{\rho}})$ and ${f}_{\rm s}({\bm{\rho}})$ are piecewise quadratic, and their gradients are Lipschitz continuous. Hence, $\nabla_{\bm{\rho}}\alpha({\bm{\rho}})$ and $\nabla_{\bm{\rho}}\alpha_{\rm s}({\bm{\rho}})$ are Lipschitz continuous over $\Omega$, and the same applies to $\nabla_{\bm{\rho}}L({\bm{\rho}})$ and $\nabla_{\bm{\rho}}L_{\rm s}({\bm{\rho}})$.
Additionally, since ${\bm{\Theta}}$ is a constant matrix with $\|{\bm{\Theta}}\|_{2}=1$, and $\nabla_{\bm{\rho}}f({\bm{\rho}})$ is Lipschitz continuous, the composition $\nabla\mathcal{L}({\bm{\varpi}})={\bm{\Theta}}\nabla L({\bm{\rho}})$ is also Lipschitz continuous. In other words, there exists a constant $C>0$ such that
$\small\|\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{1})-\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{2})\|_{2}\leq C\|{\bm{\varpi}}_{1}-{\bm{\varpi}}_{2}\|_{2},\forall{\bm{\varpi}}_{1},{\bm{\varpi}}_{2}\in\Omega_{\bm{\varpi}}$
@@
VI-B Complexity Analysis
We now qualitatively analyze and compare the computational complexity of the proposed algorithms and their benchamrks.
We observe that the increased computations of the PP-MCG-ILS algorithm, compared to the PP-MSD-ILS algorithm, primarily stems from the computation of $\varsigma_{i}^{\rm b}$ and $\varsigma_{i}^{\rm HS}$ at each iteration, which involves approximately $4N$ multiplications and 2 divisions. In both algorithms, the dominant computational burden lies in evaluating the objective function during the ILS and computing the gradients for direction updates, which are inherently tied to the CRLB expressions. Therefore, a more meaningful comparison of the overall computational cost focuses on the number of ILS steps performed per iteration. This number is influenced by the step size threshold, given by the right-hand side of (C), the initial step size $\upsilon_{0,i}$ in (48), and the scaling factor. At a given iterate ${\bm{\rho}}_{i}$, a larger search direction norm $\|{\bm{d}}_{i}\|$ generally leads to a smaller $\upsilon_{0,i}$ and a reduced step size threshold. As a consequence, the number of ILS steps per iteration is expected to remain comparable between the PP-MCG-ILS and PP-MSD-ILS algorithms. Accordingly, the primary metrics for evaluating the relative performance of these two algorithms are their convergence rate and the quality of their steady-state solutions.
We now qualitatively compare the computational complexity between algorithms employing ILS and those without ILS. As previously discussed, the number of ILS steps significantly impacts the number of evaluations of the objective function. Thus, the computational complexity for step size selection of the benchmark algorithms can be considerably lower at each iteration compared to the PP-MCG-ILS and PP-MSD-ILS algorithms. However, all four algorithms have similar complexity in direction update due to the inherently intensive gradient calculations involved. Moreover, the overall computational complexity heavily depends on the convergence rates, or equivalently, on the total number of iterations required for convergence. As will be demonstrated in the subsequent simulation results, the proposed PP-MCG-ILS algorithm exhibits a notably higher convergence rate compared to its counterparts, with or without ILS. This superior convergence performance significantly reduces the overall computational effort, particularly in the sense that it benefits demanding gradient computations.
VII Simulation Results
In this section, we conduct numerical simulations to evaluate the proposed PP-MCG-ILS, PP-MSD-ILS and P-NCG-ILS algorithms for PA optimization. All APs are assumed to be stationary. We employ the orthogonal chirp division multiplexing (OCDM) [44] waveform with a single Gaussian pulse shaping. The carrier frequency is $3$GHz. The lowpass equivalent is formulated by
@@ VII Simulation Results
VII-A CRLB Validation
We first verify the CRLB in comparison with the MLE results. To maintain consistency with existing works [28, 38, 29], we use the signal energy-to-noise ratio (SENR), defined as $\tilde{\delta}_{\rm cn}\triangleq P/\sigma_{\rm cn}^{2}$, instead of the SCNR $\delta_{\rm cn}$ in evaluating the CRLBs. Due to the high complexity of four-dimensional MLE, we consider a simple $4\times 3$ nonsymmetric multi-static radar configuration given in Fig. 2. Without loss of generality, we consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=1$ kHz.
@@ VII-A CRLB Validation
VII-B Power Allocation for ISAC
The performance of the proposed PA algorithms is evaluated by comparison with the PP-NCG and PP-NSD benchmark algorithms. To ensure communication quality, a larger number of APs are designated for transmitting ISAC signals, while a smaller subset is used for receiving target-reflected signals. Without loss of generality, we consider a CF MIMO-ISAC system comprising 10 transmitters and 2 receivers, as illustrated in Fig. 2.
To guarantee the CRLB effectiveness, we assume the minimum PA coefficient vector is ${\bm{\rho}}_{\min}=0.01\times{\bm{1}}_{N}$. The PA budget is ${\bm{\rho}}_{\max}=0.3\times{\bm{1}}_{N}$. The SENR is set to be $-10$ dB. Within the coherence time for RCS, a typical modulus-squared RCS matrix is given by $\begin{bmatrix}0.37&0.70&1.38&0.65&2.40&0.18&0.25&0.82&0.14&0.35\\ 0.40\!&\!2.05&1.14&0.42&0.03&0.11&2.02&1.65&1.24&3.65\end{bmatrix}^{\rm T}$. Similarly, the channel gains between the user and the transmitters are assumed to be ${\bm{g}}=[2.11,12.57,5.63,0.75,0.61,1.75,0.20,2.34,14.79,9.68]^{\rm T}$. We consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=10$ kHz. Under the given ISAC system setup and waveform configuration, the CRLB thresholds for position and velocity are set to $[250,0.13]$.
@@ VII-B Power Allocation for ISAC
The PA vector is initialized as the uniform PA vector ${\bm{\rho}_{1}}={\bm{1}}_{N}/N$, unless noted otherwise. The initial penalty factor is set to be $\mu_{1}=10^{4}$. We hereafter consider the ILS controlling factor $\epsilon_{L}=1e-3$, the initial step size controlling factor $\epsilon_{\upsilon}=0.9$, the penalty factor scaling coefficient $\phi=10$, and the penalty threshold $\epsilon_{\mu}=10^{-3}$, unless noted otherwise. The other simulation conditions are given in Table I.
We begin by evaluating the convergence and steady-state performance of the proposed algorithms. The SINR results corresponding to the iterative PA vectors for all four algorithms are depicted in Fig. 4. As observed, the PP-MCG-ILS algorithm exhibits the fastest convergence, followed by the PP-MSD-ILS. In contrast, the benchmark algorithms converge significantly more slowly. Furthermore, the benchmarks without ILS display noticeable numerical fluctuations, particularly under large fixed step sizes, suggesting potential instability and the risk of divergence. To ensure convergence reliability, we refrain from further increasing the step sizes in these simulations.
[…]
@@
Figure 4: SINR convergence behavior.
@@
The sequential penalty-based algorithms above leverage the optimal solution from the previous iteration loop as the initial point for the current loop under an updated penalty factor. To further explore the impact of initialization, we consider fixing the penalty factor and comparing the ISAC performance under different PA initialization strategies. In addition to the commonly used uniform PA initialization, we introduce a heuristic PA scheme, where the initial allocation is proportional to the communication channel gain, given by ${\bm{\rho}}_{1}={\bm{g}}/\sum_{n}|g_{n}|^{2}$. The results are illustrated in Figs. 7 and 8. When comparing to sequential penalty-based counterparts in Fig. 4, the PP-MCG-ILS and PP-NCG algorithms initialized with uniform PA exhibit similar convergence behavior. However, the PP-MSD-ILS and PP-NSD algorithms show degraded convergence, highlighting the superior convergence properties of CG-based algorithms over SD-based methods under identical conditions. When employing the heuristic PA initialization, all algorithms demonstrate significantly better convergence rates compared to the uniform PA case. Moreover, both PP-NCG and PP-NSD achieve noticeable performance gains relative to their sequential penalty-based counterparts, particularly PP-NCG. In contrast, the PP-MCG-ILS algorithm shows robust and superior convergence behavior, and PP-MSD-ILS exhibits a slight degradation, relative to their sequential penalty-based counterparts.
@@
We now investigate the impact of the CRLB thresholds on the performance of the proposed PP-MCG-ILS algorithm. As shown in Fig. 9, relaxing the location CRLB thresholds leads to a gradual increase in the steady-state SINR, which eventually stabilizes. This trend aligns with the behavior observed in Fig. 9, where the achieved location CRLB value closely approaches its threshold and then converge. These trends reveal the trade-off between the communication SINR and the location CRLBs. Furthermore, it is noteworthy that the velocity CRLB remains largely unaffected by variations in the location CRLB thresholds. In contrast, Fig. 10 shows that the performance of the PP-MCG-ILS algorithm exhibits negligible variation with the increased velocity CRLB threshold. The observed CRLB saturation with respect to the corresponding CRLB threshold may be an inherent outcome of the introduced power budget constraints.
@@
VII-C Power Allocation for Pure Sensing
@@
We now verify the effectiveness of the proposed P-MCG-ILS algorithm for total power minimization in pure sensing scenario with joint location and velocity estimation. The initial penalty factor is set to be $\mu_{1}=1$, $\sigma_{\upsilon}=0.2$ and the penalty threshold $\epsilon_{\mu}=10^{-6}$. The other simulation conditions are the same as the PP-MCG-ILS algorithm in Fig. 4.
@@
The iterative results for total power optimization and the corresponding CRLB performance are illustrated in Fig. 11. As seen in Fig. 11, the algorithm demonstrates rapid convergence. Under the same CRLB threshold settings, the total power consumption in the pure sensing scenario is significantly lower than that in the ISAC case. This is expected, as ISAC systems must allocate sufficient power to meet both sensing and communication requirements, particularly for maximizing communication performance. Moreover, both the location and velocity CRLBs in the pure sensing case closely align with their respective thresholds. In contrast, in the ISAC scenario (refer to Fig. 6), the velocity CRLB matches its threshold loosely. This difference arises because, in pure sensing, any enhancement in CRLB performance directly benefits from increased sensing power. However, in ISAC, PA is optimized primarily for communication performance, with CRLB constraints being satisfied, potentially loosely, within acceptable margins.
VIII Conclusions